### v1

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Metamaterials are a promising platform for a range of applications, from shock absorption to mechanical computing. These functionalities typically rely on floppy modes or mechanically frustrated loops, both of which are difficult to design. Here, we introduce a combinatorial approach that allows to create an arbitrarily large number of floppy modes and frustrated loops. The design freedom of the mode shapes enables us to easily introduce kinematic incompatibility to turn them into frustrated loops. We demonstrate that floppy modes can be sequentially buckled by using a specific instance of elastoplastic buckling, and we utilize the designability of floppy chains and frustrated loops to demonstrate matrix-vector multiplication in materia.
Introduction – Mechanical metamaterials with programmable mechanical response show promise for applications ranging from wave guiding Huber (2016) and shock absorption Liu et al. (2024), to shape morphing Dudek et al. (2025) and mechanical computing Treml et al. (2018); Yasuda et al. (2021). Two particularly useful paradigms for the design of metamaterials are (i) floppy modes, which are deformations that cost negligible elastic energy Lubensky et al. (2015); Haghpanah et al. (2016); Meng et al. (2022); Bertoldi et al. (2017); Coulais et al. (2016); Dieleman et al. (2020); Bossart et al. (2021); van Mastrigt et al. (2022, 2023); Sirote-Katz et al. (2024a); Louvet et al. (2024a); Deng et al. (2020); Czajkowski et al. (2022); and (ii) frustrated loops Kang et al. (2014); Meeussen et al. (2020a, b); Pisanty et al. (2021); Guo et al. (2023); Sirote-Katz et al. (2024b, c), where these low-energy deformations are geometrically impossible. Floppy modes and frustrated loops can be used to channel deformations and stresses efficiently, yet designing metamaterials with multiple, precisely controlled modes or loops remains challenging.
To address this challenge, we open up the design space of combinatorial metamaterials, by introducing an approach that allows much freedom in setting the number and shapes of floppy modes and frustrated loops. We demonstrate how to apply this for sequential response and mechanical computing. Our method for designing multifunctional mechanical metamaterials uses mixtures of unimodal and bimodal triangular building blocks that are fully described by a spin model characterizing spatially complex floppy modes, which we can in turn design. Unlike other systems, which mainly rely on trial and error Bossart et al. (2021) or on computational methods van Mastrigt et al. (2022) for their design process, our work provides a framework for systematically designing complex mechanical metamaterials, bridging abstract principles and practical applications in soft programmable materials and adaptive architecture using a simple intuitive model.
Figure 1: Floppy modes in ideal metamaterials. a, Triangular blocks made of perfect hinges and rigid bonds. Spins (green) describe the direction of displacement of edge nodes. Bonds (brown) constrain spins to move in an alternating manner. One internal bond in Block $T_{1}$ constrains two spins, such that the third spin is independent, resulting in two floppy modes. Block $T_{2}$ has two internal bonds, thus all spins displace together, resulting in one floppy mode. b, Theoretical design (left) and experimental construction (right) of metamaterial made from $T_{1}$ and $T_{2}$ blocks. Chains of connected spins that constrain each other’s motion are individually colored; open chains (light green, yellow, and orange) represent nodes moving together in a floppy mode, and so do chains that contain closed loops of even length (dark green). Chains with an odd loop (red) are mechanically frustrated and thus rigid. c, Deformations showcasing the four floppy modes of the system in b. d, Number of floppy modes $F$ as function of the number $N_{1}$ of $T_{1}$ blocks for an extended lattice of $N=210$ blocks, where both $F$ and $N_{1}$ are scaled by $N$. The achieved values of $F$ span almost the entire range between the lower and upper theoretical bounds. The average number of floppy modes in randomly generated systems exhibits small fluctuations and is closer to the lower bound.
Building Blocks and Spin Model – We consider two-dimensional mechanical metamaterials composed of triangular building blocks, which are theoretically constructed from rigid bonds connected at freely rotating hinges. Each block consists of three corner nodes and three edge nodes. The nodes are connected by six rigid perimeter bonds. In addition, $T_{1}$ triangles have one internal bond connecting two adjacent edge nodes, while $T_{2}$ triangles have two such internal bonds (Fig. 1a). For small floppy deformations, the corner nodes remain stationary while edge nodes move perpendicular to their respective edges. This key insight allows us to introduce a spin model to describe the system’s deformations. In this model, we assign a spin-like variable to each edge node, describing whether it displaces in to the triangle or out of it. Effective antiferromagnetic interactions between spins are introduced by internal bonds, since they constrain connected edge nodes to move in alternating directions, with respect to the triangle.
Design of Floppy Modes – Using our spin model, we can design metamaterials with varying numbers and shapes of floppy modes. By connecting spins we create chains of nodes (Fig. 1b) that can move together, independently from other chains in the system (Fig. 1c). The mutual direction of motion of the nodes is determined by the internal bonds connecting them along the chain. This demonstrates our ability to create chains with complex, predefined shapes, and deformations. Furthermore, an inherent property caused due to the antiferromagnetic interaction between spins is that a closed loop with an even number of such bonds allows the corresponding spins to move together (dark green chain in Fig. 1b,c), while a loop with an odd length renders the chain frustrated and rigidifies the connected spins (red chain in Fig. 1b).
To validate our approach and the spin model, we explicitly write the rigidity matrix for the actual geometries that we consider, we identify from the matrix how many floppy modes the system has, and then verify that the theoretically predicted floppy modes from the model indeed give zero when multiplied by the rigidity matrix. This shows that we can use the deformations given by the spin model as the base of all possible non-trivial deformations of the system. We also construct a physical model using LEGO® technic beams and axles, which closely approximates our theoretical system of flexible hinges and rigid bonds. We successfully actuated the floppy modes by displacing nodes, which validates our design methodology and shows that these modes are, in fact, independent zero-energy deformation modes (Fig. 1c). How many such floppy modes are there in this class of metamaterials?
Number of Modes – The number of floppy modes is essentially the number of independent degrees of freedom in the system. Each spin is a degree of freedom, and each bond is a constraint that reduces the number of independent degrees of freedom. However, in each closed loop one bond is redundant and does not add a constraint, and chains with an odd loop are rigid and do not possess a degree of freedom. A lattice of $N$ triangles and of perimeter $P$ contains $\frac{3N}{2}+\frac{P}{2}$ edge nodes, or spins. Denoting the numbers of $T_{1}$ and $T_{2}$ triangles by $N_{1}$ and $N_{2}$, respectively, the number of internal bonds is $N_{1}+2N_{2}$. Using $N=N_{1}+N_{2}$, we obtain the following exact expression for the number of floppy modes,
$F=N_{1}-\frac{N}{2}+\frac{P}{2}+L-R,$ (1)
where $L$ is the number of closed loops of internal bonds, and $R$ is the number of rigid chains.
For given lattice size and shape that set $N$ and $P$, and for given values $N_{1}$ and $N_{2}$ of $T_{1}$ and $T_{2}$ blocks, the number of floppy modes can take a wide range (see Fig. 1d), depending on how the blocks are arranged to form chains and loops, thus changing the value of $L-R$. The metamaterial design can lead to any number of modes between the bounds,
$F\geq\begin{cases}0,&N_{1}\leq\frac{N-P}{2}\\ N_{1}-\frac{N-P}{2},&N_{1}\geq\frac{N-P}{2}\end{cases},$
$F\leq\begin{cases}N_{1}+1,&N_{1}\leq\frac{3P}{2}-3\\ \frac{2N_{1}}{3}+\frac{P}{2},&N_{1}\geq\frac{3P}{2}-3\end{cases}.$ (2)
Thus, if most triangles are $T_{2}$ $\left({\rm namely,}\ N_{1}\lesssim\frac{N}{2}\right)$, there is a way to arrange the triangles such that the metamaterial will be rigid, without any floppy modes, $F=0$. For any ratio of $T_{1}$ and $T_{2}$ triangles, there are always designs that give an extensive number of floppy modes $F\propto N$, and if most blocks are $T_{1}$ $\left(N_{1}\gtrsim\frac{N}{2}\right)$, this number is necessarily extensive. Finally, we observe that by randomly positioning and orienting the blocks, the number of floppy modes is closer to the lower bound (Fig. 1d).
Figure 2: Linked floppy modes and rigid chains in three dimensions. a, In a single layer, two crossing chains (red and yellow) share an edge node (black star) that couples them. b, Connecting parallel layers allows chains to bypass each other. c, Three dimensional metamaterials enable knotted topologies, like catenated floppy chains, without contact between them. For clarity, the vertical connectors at all corner nodes are not shown. d, LEGO® realization of the linked loops and their individual actuation. Moving the solid black vertical connector to its dotted position rigidifies the yellow loop without changing the intra-layer structure.
Mode Crossing – Although we can design the shapes of multiple floppy modes, these modes cannot cross each other. For that, they would have to share a node at the crossing point. That node’s motion would cause the simultaneous actuation of the two modes, and they would thus, in fact, be a single mode (Fig. 2a). To allow modes to cross, we generalize our approach to a three-dimensional layered system where modes can bypass each other via the third dimension (Fig. 2b). We couple adjacent layers with vertical connectors and allow only planar displacements within each layer. By connecting corner nodes across layers, we remove shearing and twisting.
Connecting specific edge nodes we can design floppy modes of arbitrary topology, like the catenated floppy loops shown in Fig. 2c that may be individually actuated since they do not touch each other (Fig. 2d). Furthermore, chains that in individual layers are floppy may now form rigid loops that traverse multiple layers in a manner that is sensitive to the positions of the vertical connectors. For instance, moving the black connector in Fig. 2c to its dotted position changes the parity of the yellow loop from even and thus floppy to odd and thus rigid, without changing the structure of any of the layers.
By extending our design methodology to layered lattices, we significantly expand the range of achievable functionalities and topologies in our mechanical metamaterials, bridging the gap between planar designs and fully three-dimensional structures.
Figure 3: Sequential actuation of floppy modes in flexible metamaterials. a, Theoretical model and corresponding metamaterial geometry, in which rigid bonds are replaced with diamonds. b, Metamaterial with three floppy chains (red) of varying length and curvature, separated by disconnected rigid chains (blue). c, 3D-printed metamaterial. d, Elastic metamaterial at different stages of compression. f Compression of plastic metamaterial of the same design. h, Compression of plastic metamaterial with three identical floppy chains. Color-coded arrows in d, f, and h indicate the horizontal displacement. e, g, i, Corresponding vertical force, $F$ and average horizontal displacement, $u_{1}$ of each floppy chain vs. global compressing stroke $u/H$. Numbers indicate the differently shaped chains, and $B$, $M$, and $T$ denote the identically shaped bottom, middle, and top floppy chains in h. The sign of the slope ($S$) of force vs compression is detected (dashed lines).
Buckling of Floppy Modes – Floppy modes and frustrated loops can be used as design tools; we go beyond the local actuation of individual modes, and exploit the buckling instability as a mechanism to actuate floppy modes by uniform global compression of the metamaterial. This allows us to tailor the system’s force-displacement curve. Unlike existing metamaterials based on buckling, which have a single mode Bertoldi et al. (2010), a large number of modes that are located along straight lines Lubensky et al. (2015); Bossart et al. (2021); van Mastrigt et al. (2022); Liu et al. (2024, 2025), or limited modes in a hierarchical design Coulais et al. (2018a); Wu and Pasini (2024), our modes can be shaped into tortuous chains or loops, and the buckling onset of those modes is determined by the length of the chains and by the parity of the loops.
We demonstrate the actuation of floppy modes in a 3D-printed metamaterial composed of diamond-shaped blocks (Fig. 3a). Here, the thin connectors at the ends of the diamonds allow a relatively flexible rotation compared to the substantially greater resistance to deforming the individual diamonds. We first design three roughly horizontal floppy chains of different lengths and curvatures across the metamaterial and separate them by rigid-loop layers (Fig. 3b). To trigger the buckling of the floppy modes, we apply a uniform vertical compression to the metamaterial. We study elastic buckling by 3D printing with a rubber-like polymer (Fig. 3c). Under compression, the three floppy modes in the elastic metamaterial buckle almost simultaneously (Fig. 3d,e) because of the close buckling loads of each floppy mode and the positive stiffness at the onset of the elastic buckling (Fig. 3e).
Sequential Yield Buckling – We separate the excitation of the different modes by exploiting yield buckling Liu et al. (2024), which is rooted in the symbiotic occurrence of plastic yielding and buckling. To combine plasticity with the floppy modes, we 3D print a metamaterial with the same geometry but made of an elastoplastic material. Under quasistatic compression, the straight mode of lowest energy buckles first, followed by a negative stiffness, until all the diamond components in the first buckled layer reach self-contact (Fig. 3f left). At this point, the metamaterial stiffens until the critical buckling load of the second-length mode is reached (Fig. 3f middle). Such sequential buckling repeats until the last mode with the longest length buckles and reaches contact. Due to this sequential buckling, the force curve exhibits a wiggly increasing plateau with tunable local maxima (Fig. 3g). Thus, our designed floppy modes can be used to tailor the nonlinear mechanical response. A metamaterial in which all floppy chains have the same length and curvature also exhibits sequential buckling with negative stiffness at the onset of yield buckling (Fig. 3h,i). In this case, the buckling order of the modes is determined by imperfections and boundary conditions. These findings could be further leveraged to optimize vibration Dykstra et al. (2023); Hu et al. (2023); Meng et al. (2022) and shock absorption Liu et al. (2024, 2025); Shan et al. (2015); Tancogne-Dejean et al. (2016) performances.
Matrix-Vector Multiplication – Computing in materia is an emerging direction Momeni et al. (2024); Yasuda et al. (2021), studied in a myriad of physical platforms, such as cross-bar arrays Xia and Yang (2019), spintronic devices Ahn (2020), electromagnetic Estep et al. (2014); Silva et al. (2014) and phononic coupled resonators Okamoto et al. (2013), and microfluidics Ahrar et al. (2023). For these, the ability to perform algebraic operations is paramount, and therefore a crucial challenge is how to achieve matrix manipulations. Recent work has shown that floppy modes can be used to achieve matrix-vector multiplication in mechanical systems Louvet et al. (2024b). Here, we demonstrate that frustrated loops are an efficient way to perform algebraic operations with multiple inputs and outputs.
Figure 4: Matrix-vector multiplication a, experimental setup wherein the six corner nodes are fixed in place. b, experimental results (dots) and theoretical predictions with $\alpha=0.8$ (solid line) of output under a single input from the top, $I_{\text{T}}$ or the bottom, $I_{\text{B}}$. Colors correspond to the outputs marked in panel a. c, experimental measurements (solid lines) and predictions (dashed lines) under two simultaneous inputs. d, three-layer metamaterial. e, corresponding 3D structure. f, experiment results of inputs in the middle layer, and outputs in the left and right layers. The solid lines are the measurements and the dashed lines are predictions.
The minimal metamaterial with a frustrated loop consists of six $T_{2}$ blocks (Fig. 4a). To enable manual actuation, we 3D print the diamond-shaped beams in this metamaterial with a flexible material and print the ligaments with a softer material. We fix the hexagon’s six corner nodes and locally actuate the top and bottom edge nodes, as input displacements $I_{\rm T}$ and $I_{\rm B}$. In the ideal model of rigid bonds, realized above using LEGO® beams, a chain containing an odd loop is rigid, and may not deform. However, the 3D-printed metamaterial has some flexibility, leading to non-trivial output displacements $O_{\rm TL}$, $O_{\rm TR}$, $O_{\rm BL}$, and $O_{\rm BR}$.
Actuating each of the inputs separately, we observe linear behavior of each of the outputs (Fig. 4b). Thus, this frustrated loop relates the input and output vectors,
$\vec{I}=\begin{pmatrix}I_{\rm T}\\ I_{\rm B}\end{pmatrix},\quad\vec{O}=\begin{pmatrix}O_{\rm TL}\\ O_{\rm TR}\\ O_{\rm BL}\\ O_{\rm BR}\end{pmatrix}$ (3)
via the matrix-vector product,
$\vec{O}=A\vec{I}.$ (4)
Each internal bond’s effective antiferromagnetic interaction flips the displacement’s direction. Due to the odd length of frustrated loops, the transmission of displacement from the input node along the loop’s two branches leads to displacement in opposite directions at the output node. However, beyond these flips in direction, the elastic nature of the 3D-printed metamaterial causes a decay in the displacement’s magnitudes Coulais et al. (2018b), which we theoretically describe by each beam decreasing the displacement by some fixed factor $\alpha$. Tracking the two paths from each input to each output node in the hexagonal structure of Fig. 4a, we obtain the transmission matrix,
$A$ $=\begin{pmatrix}-\alpha^{4}+\alpha^{7}&-\alpha^{2}+\alpha^{7}&\alpha^{4}-\alpha^{7}&\alpha^{4}-\alpha^{7}\\ -\alpha^{3}+\alpha^{8}&\alpha^{4}-\alpha^{5}&\alpha^{3}-\alpha^{8}&-\alpha^{4}+\alpha^{7}\end{pmatrix},$ (5)
which is consistent with the experimental measurement (Fig. 4b). Simultaneously applying two inputs in the same direction and then in opposite directions, we see in Fig. 4c that superposition holds, such that the output displacements are indeed given by Eq. (4).
We have shown how each frustrated loop may receive two inputs. Multiplying larger vectors is possible by feeding the output from one loop as an input to another loop. However, this requires transmission lines to cross, which is not possible in 2D metamaterials. We overcome this by coupling adjacent layers in a 3D metamaterial (Fig. 4d,e). Individually actuating each of the input nodes in the middle layer, and measuring the response of each of the output nodes in the left and right layers, we obtain all elements of the transmission matrix $A$. We then test our approach (Fig. 4f) by simultaneously applying different combinations of several inputs, and observe that the output displacements are well described by Eq. (4).
We introduced a combinatorial approach using unimodal and bimodal triangular building blocks to create metamaterials with arbitrary shapes and a large number of floppy modes and frustrated loops. We demonstrated that curved floppy modes together with yield buckling can be used to achieve sequential deformations, and that frustrated loops can be exploited to achieve matrix-vector multiplication. Our findings enrich the toolbox of sequential metamaterial design and lead to various applications in shock and vibration damping, shape changing, and mechanical computing. Our work opens questions like: can we use plasticity and fatigue to learn in combinatorial metamaterials Du et al. (2025)? How to integrate matrix-vector multiplication and nonlinear response like buckling for the function of logic gates Guo et al. (2023)? How can floppy modes and frustrated loops be used for locomotion in active metamaterials Veenstra et al. (2025)?
Acknowledgments – We thank Chaviva Sirote-Katz, Dor Shohat and Marc Serra-Garcia for inspiring discussions and Bat-El Pinchasik and Yaara Shokef for technical assistance. T.A.S. and Y.S. thank the Institute of Physics at the University of Amsterdam for its hospitality. W.L. and C.C. acknowledge funding from the European Research Council under grant agreement 852587 and the Netherlands Organisation for Scientific Research under grant agreement NWO TTW 17883. C.C. acknowledges funding from the Netherlands Organisation for Scientific Research under grant agreement NWO under grant agreement VI.Vidi 2131313. This research was supported in part by the Israel Science Foundation Grant No. 1899/20.
I Appendix
I.1 Bounds on Number of Floppy Modes
To maximize the number $F$ of floppy mode, we minimize the number $R$ of rigid chains and maximize the number $L$ of loops. The minimal value of $R$ is zero. We now derive two upper bounds on $L$. The first stems from the fact that each loop surrounds at least one internal corner node, thus $L\leq\frac{N+2-P}{2}$, which leads to $F\leq N_{1}+1$.
For the second bound, we start with a lattice of $N_{2}=0$ and all triangles as $T_{1}$; In this limit, the maximal number of loops givs $L\leq\frac{N}{6}$, by arranging the loops in a honeycomb formation. Now, as we gradually switch $T_{1}$ blocks to $T_{2}$, the most efficient way to create new loops is that two loops share each $T_{2}$ block. Each new loop needs six bonds, thus $L\leq\frac{N}{6}+\frac{N_{2}}{3}$, and $F\leq\frac{2N_{1}}{3}+\frac{P}{2}$. The crossover between these two bounds occurs at $N_{1}=\frac{3P}{2}-3$, and we obtain the upper bound given in Eq. (2).
For a given lattice of size and shape, the system becomes more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with $E-1$ bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in Eq. (1) we obtain the lower bound given in Eq. (2).
In the large system size limit, $N\gg 1$, since $P\ll N$, the lower bound is approximated by
$F\geq\begin{cases}0,&N_{1}\leq\frac{N}{2}\\ N_{1}-\frac{N}{2},&N_{1}\geq\frac{N}{2},\end{cases}$ (6)
and the upper bound by $F\leq\frac{2N_{1}}{3}$. As system size increases, the average number of floppy modes in randomly generated lattices collapses to a single curve, see Fig. 5.
Figure 5: Average number of floppy modes in randomly generated lattices. a, Results for different system sizes normalized by $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, such that the lower bound will coincide. b, Normalized umber of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
I.2 Matrix-Vector Multiplication with Multi-Layer Rigid Loops
We have demonstrated the matrix-vector multiplication in a single-layer metamaterial. Here, we further show a larger matrix-vector multiplication in a three-layer metamaterial of four inputs in the middle layer and 12 outputs in the side layers. These frustrated loops similarly relate the input and output vectors. We experimentally apply a single input and measure the transmission matrix $A$ from the slopes of the input-output curves (Fig. 6). Similarly to the single-layer case, we predict the transmission by assuming decay factor $\alpha=0.8$ within each layer, decay factor $\beta_{1}=0.4$ when the input transmits from one layer to another layer, and $\beta_{2}=1$ when two layers have the same input, like the output in the left layer $O_{\rm T}=I_{\rm T}$ (Fig. 4d). We find good agreement between the theoretical predictions and the experimental measurements (Fig. 6).
Figure 6: Experimental outputs of the three-layer metamaterial with single inputs. a, Left layer. b, Right layer. Inputs are applied to the middle layer. Solid lines are the theoretical predictions.
Figure 7: 3D printed metamaterials and experimental method. a, A 3D printed metamaterial for buckling with elastoplastic material (Nylon). b, A 3D printed metamaterial for buckling with rubber-like polymer (Polyject Agilus + Vero). c, A metamaterial with tracking markers at the connected joints. d, A single-layer metamaterial for local actuation is printed with two materials. Purple ligaments and red diamonds are printed with soft material (Agilus) and rigid material (Vero), respectively. e, Experimental setup of a three-layer metamaterial with local actuation under filming of two cameras. The picture in the top-right corner is the recorded image from the right camera
I.3 Experimental methods
The metamaterials with floppy modes designed in this work consist of a plurality of triangular unit cells, wherein the rigid bonds and rotating nodes from the theoretical model are replaced with solid diamond blocks of length $\ell$ and ligaments of thickness $t$, respectively. For the metamaterials with buckling floppy modes, we 3D printed them with a single material of $\ell=10.3$ mm, $t=0.4$ mm, and $d=3$ mm. For the metamaterials with elastoplastic buckling (Fig. 7a), we used an elastoplastic material, nylon, with the fused deposition modelling (FDM) technology (Ultimaker 2S). For the metamaterials with elastic buckling (Fig. 7b), we 3D printed them with a rubber-like polymer material (Stratasys, Aligus+Vero). The thicknesses of the metamaterials are set to $T=24.5$ mm for elastoplastic metamaterials and $T=23.2$ mm for elastic metamaterials to prevent out-of-plane buckling. We then marked the ligaments on one side of the metamaterials with white dots for image tracking (Fig. 7c). We used Image-J software to track the position of the white dots of each frame in the video. For the metamaterials with local actuation, we changed the design of ligaments from a rectangular section to a section with inserted tapered rigid parts (Fig. 7d). We then printed such ligaments with a soft material (Aligus) and the diamonds with a rigid material (Vero). Such ligament design can significantly increase resistance to shear and tension while offering only a negligible resistance to bending  Bossart et al. (2021). For the single-layer metamaterial (Fig. 4b), we set the thickness of the sample to 6 mm. For the three-layer metamaterials with multiple inputs and outputs, we printed each layer separately and assembled them with two rigid frames. The middle layer has a thickness of 6 mm and the left and right layers have a thickness of 3.6 mm. We set a small distance of 0.5 mm between each layer to avoid friction during actuation. We then fix the sample on the table. We marked the ligaments at the six edges of the hexagon layer with black dots (Fig. 7e top-right corner). We then manually pull and push the rigid bars of the middle layer and record the movement of the 12 dots from the left and right layers with two Basler cameras (Fig. 7e).