### v2

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Metamaterials are a promising platform for a range of applications, from shock absorption to mechanical computing. These functionalities typically rely on floppy modes or mechanically frustrated loops, both of which are difficult to design. We introduce a combinatorial approach that allows to create an arbitrarily large number of floppy modes and frustrated loops. The design freedom of the mode shapes enables us to easily introduce kinematic incompatibility to turn them into frustrated loops. We demonstrate that floppy modes can be sequentially buckled by using a specific instance of elastoplastic buckling, and we utilize our combinatorial floppy chains and frustrated loops to achieve matrix-vector multiplication in materia. Our findings bring about new principles for the design and the use of floppiness and geometric frustration in soft matter and metamaterials.
Introduction – In soft matter, mechanical properties are often rooted in floppy modes and geometrically frustrated states of self-stress. Examples range from protein allostery [1], granular packings [2, 3] and colloidal glasses [4, 5] to amorphous solids [6, 7, 8], biopolymer networks [9, 10, 11] and mechanical metamaterials [12, 13]. Mechanical metamaterials are a platform of choice to study floppy and frustrated modes [14]. Once one understands those modes, one can in turn harness them to achieve on-demand unusual properties [15]. Prime examples are topological wave guiding and stress focusing [16, 17, 18, 19, 20, 21, 22, 23], vibration and shock absorption [24, 25, 26], shape morphing [27, 28, 29, 30] and mechanical computing [31, 32, 33, 34, 35, 36].
Floppy modes are deformations that cost negligible elastic energy [37, 38, 39, 40, 41, 42, 43, 44]. Frustrated stressed states are often configured in loops [45, 46, 47, 48, 49, 50, 51], where these low-energy deformations are geometrically impossible. Both floppy modes and frustrated loops can be used to channel deformations and stresses efficiently, yet designing metamaterials with multiple, precisely controlled modes or loops remains a challenge. To address this problem, we open up the design space of combinatorial metamaterials, by introducing a theoretical approach that allows much freedom in setting the number and shapes of floppy modes and frustrated loops. We leverage this approach to create metamaterials that exhibit advanced sequential response upon uniaxial compression and mechanical computing in the form of matrix-vector multiplication. Thus, our work bridges abstract principles and practical applications in soft programmable materials and adaptive architectures.
Figure 1: Floppy modes in metamaterials with perfect hinges and rigid bonds. a, Triangular blocks with spins (green) describing the direction of displacement of edge nodes, and bonds (brown) constraining spins to move in alternating manner. One internal bond in Block $T_{1}$ constrains two spins, thus the third is independent, resulting in two floppy modes. Block $T_{2}$ has two internal bonds, thus all spins displace together in one floppy mode. b, Metamaterial made of $T_{1}$ and $T_{2}$ blocks. Chains of connected spins constraining each other’s motion are individually colored; open chains (light green, yellow, orange) represent nodes moving together in a floppy mode, and so do chains that contain closed loops of even length (dark green). Chains with odd loop (red) are mechanically frustrated and hence rigid. c, Experimental demonstration of the four floppy modes of the system in b. d, Normalized number $F/N$ of floppy modes vs. normalized number $N_{1}/N$ of $T_{1}$ blocks for a lattice of $N=210$ blocks. The values of $F$ span almost the entire range between the lower and upper theoretical bounds. The average number of floppy modes in randomly generated systems exhibits small fluctuations and is closer to the lower bound.
Building Blocks and Spin Model – We consider two-dimensional mechanical metamaterials composed of triangular building blocks, which are constructed from rigid bonds connected at freely rotating hinges. Each block consists of three corner nodes and three edge nodes. The nodes are connected by six rigid perimeter bonds. In addition, $T_{1}$ triangles have one internal bond connecting two adjacent edge nodes, while $T_{2}$ triangles have two such internal bonds (Fig. 1a). For floppy modes, in the small deformation limit, the corner nodes remain stationary while edge nodes move perpendicular to their respective edges. This key insight allows us to introduce a spin model to describe the system’s deformations. In this model, we assign a spin-like variable to each edge node, describing whether it displaces in to the triangle or out of it. Effective antiferromagnetic interactions between spins are introduced by internal bonds, since they constrain connected edge nodes to move in alternating directions with respect to the triangle.
Design of Floppy Modes – Using our spin model, we can design metamaterials with varying numbers and shapes of floppy modes. By connecting spins, we create chains of nodes (Fig. 1b) that can move together, independently from other chains in the system (Fig. 1c). The mutual direction of motion of the nodes is determined by the internal bonds connecting them along the chain. This demonstrates our ability to create chains with complex, predefined shapes, and deformations. Furthermore, an inherent property caused due to the antiferromagnetic interaction between spins is that a closed loop with an even number of such bonds allows the corresponding spins to move together (Fig. 1b,c, dark green chain), while a loop with an odd length renders the chain frustrated and rigidifies the connected spins (Fig. 1b, red chain).
To validate our approach and the spin model, we explicitly write the rigidity matrix for the actual geometries that we consider, we identify from the matrix how many floppy modes the system has, and then verify that the theoretically-predicted floppy modes from the model indeed give zero when multiplied by the rigidity matrix. We construct a physical model using LEGO® beams and axles, which closely approximates our theoretical system of flexible hinges and rigid bonds. We successfully actuate the floppy modes by displacing nodes, which validates our design methodology and shows that these modes are, in fact, independent zero-energy deformation modes (Fig. 1c and Video 1).
Number of Modes – On top of designing the spatial form of floppy modes, the spin model allows to quantify their number in such metamaterials; The number of floppy modes is the number of independent degrees of freedom in the system. Each spin is a degree of freedom, and each bond is a constraint that reduces the number of independent degrees of freedom. However, a bond which closes a loop is redundant and does not add a constraint, while chains with an odd loop are rigid and do not possess a degree of freedom. A lattice of $N$ triangles and of perimeter $P$ contains $\frac{3N}{2}+\frac{P}{2}$ edge nodes, or spins. Denoting the numbers of $T_{1}$ and $T_{2}$ triangles by $N_{1}$ and $N_{2}$, respectively, the number of internal bonds is $N_{1}+2N_{2}$. Using $N=N_{1}+N_{2}$, we obtain an exact expression for the number of floppy modes, $F=N_{1}-\frac{N}{2}+\frac{P}{2}+L-R$, where $L$ is the number of closed loops of internal bonds, and $R$ is the number of rigid chains.
For given lattice size and shape that set $N$ and $P$, and for given values $N_{1}$ and $N_{2}$ of $T_{1}$ and $T_{2}$ blocks, the number of floppy modes can take a wide range (Fig. 1d), depending on how the blocks are arranged to form chains and loops, thus changing the value of $L-R$. The metamaterial design can lead to any number of modes between the bounds,
$F\geq\begin{cases}0,&N_{1}\leq\frac{N-P}{2}\\ N_{1}-\frac{N-P}{2},&N_{1}\geq\frac{N-P}{2}\end{cases},$
$F\leq\begin{cases}N_{1}+1,&N_{1}\leq\frac{3P}{2}-3\\ \frac{2N_{1}}{3}+\frac{P}{2},&N_{1}\geq\frac{3P}{2}-3\end{cases}.$ (1)
If most triangles are $T_{2}$ $\left({\rm namely,}\ N_{1}\lesssim\frac{N}{2}\right)$, the triangles can be arranged such that the metamaterial will be rigid, without any floppy modes, $F=0$. For any ratio of $T_{1}$ and $T_{2}$ triangles, there are always designs that give an extensive number of floppy modes $F\propto N$, and if most blocks are $T_{1}$ $\left(N_{1}\gtrsim\frac{N}{2}\right)$, this number is necessarily extensive. Finally, we observe that by randomly positioning and orienting the blocks, the number of floppy modes is closer to the lower bound (Fig. 1d).
Figure 2: Linked floppy modes and three-dimensional rigid chains. a, In the plane, two crossing chains (red, yellow) share an edge node (black star) that couples them. b, Connecting parallel layers allows chains to bypass each other. c, Three-dimensional metamaterials enable knotted topologies, like catenated floppy chains, without contact between them. For clarity, the vertical connectors at all corner nodes are not shown. Moving the solid black vertical connector to its dotted position rigidifies the yellow loop without changing the intra-layer structure. d, LEGO® realization of linked loops and their individual actuation.
Mode Crossing – Although we can design the shapes of multiple floppy modes, these modes cannot cross each other. For that, they would have to share a node at the crossing point. That node’s motion would cause the simultaneous actuation of the two modes, and they would thus, in fact, be a single mode (Fig. 2a). To allow modes to cross, we generalize our approach to a three-dimensional layered system where modes can bypass each other via the third dimension (Fig. 2b). We couple adjacent layers with vertical connectors and allow only planar displacements within each layer. By connecting corner nodes across layers, we remove shearing and twisting.
Connecting specific edge nodes, we can design floppy modes of arbitrary topology, like catenated floppy loops (Fig. 2c) that may be individually actuated since they do not touch each other (Fig. 2d and Video 2). Furthermore, chains that in individual layers are floppy may now form rigid loops that traverse multiple layers in a manner that is sensitive to the positions of the vertical connectors. For instance, moving the black connector in Fig. 2c to its dotted position changes the parity of the yellow loop from even and thus floppy to odd and thus rigid, without changing the structure of any of the layers.
By extending our design methodology to layered lattices, we significantly expand the range of achievable functionalities and topologies in our mechanical metamaterials and (literally) bridge the gap between planar designs and fully three-dimensional structures. In the remainder of the paper, we utilize (i) floppy modes to achieve functional responses such as sequential buckling of floppy modes with designer shapes and (ii) frustrated loops to realize matrix-vector multiplication.
Figure 3: Sequential actuation of floppy modes. a, Theoretical model and corresponding metamaterial geometry, with rigid bonds replaced by diamonds. b, Metamaterial with three floppy chains (red) of varying length, separated by disconnected rigid chains (blue). c, 3D-printed metamaterial. d, Elastic metamaterial at different stages of compression. f Compression of plastic metamaterial of the same design. h, Compression of plastic metamaterial with three identical floppy chains. Color-coded arrows in d, f, and h indicate horizontal displacement. e, g, i, Corresponding vertical force, $F$ and average horizontal displacement, $u_{1}$ of each floppy chain vs. global compressing stroke $u/H$. $B$, $M$, and $T$ denote the identically shaped bottom, middle, and top floppy chains in h. The sign of the slope ($S$) of force vs compression is detected (dashed lines).
Floppy Mode Buckling – To go beyond local actuation of individual modes, we exploit the buckling instability as a mechanism to actuate floppy modes by uniform global compression. Unlike existing metamaterials based on buckling, with one mode [52], a large number of modes along straight lines [13, 39, 41, 26, 53], or limited modes in a hierarchical design [54, 55], our modes can be shaped into tortuous chains or loops, and their buckling onset is determined by the length of the chains and by the parity of the loops (Video 3).
We demonstrate floppy mode actuation in a 3D-printed metamaterial made of diamond-shaped beams (Fig. 3a); The thin connectors at the diamond ends allow a relatively flexible rotation compared to the substantially greater resistance to deforming the individual diamonds. We design three roughly horizontal floppy chains of different lengths and separate them by rigid layers (Fig. 3b). To trigger the buckling of the floppy modes, we apply a uniform vertical compression to the metamaterial. We study elastic buckling by 3D printing with a rubber-like polymer (Fig. 3c). Under compression, the three floppy modes buckle almost simultaneously (Fig. 3d,e) because of their close buckling loads and the positive stiffness at the onset of elastic buckling (Fig. 3e).
Sequential Yield Buckling – We separate the excitation of the different modes by exploiting yield buckling [26], which is rooted in the symbiotic occurrence of plastic yielding and buckling. To combine plasticity with the floppy modes, we 3D print an elastoplastic metamaterial of the same geometry. Under quasistatic compression, the straight mode of lowest energy buckles first, followed by a negative stiffness, until all the diamonds in the first buckled layer reach self-contact (Fig. 3f left). At this point, the metamaterial stiffens until the critical buckling load of the second-length mode is reached (Fig. 3f middle). Such sequential buckling repeats until the last mode with the longest length buckles and reaches contact. Due to this sequential buckling, the force curve exhibits a wiggly increasing plateau with tunable local maxima (Fig. 3g). Thus, our designed floppy modes can be used to tailor the system’s nonlinear force-displacement response. A metamaterial with all floppy chains of the same length also exhibits sequential buckling with negative stiffness at the onset of yield buckling (Fig. 3h,i). Here, the buckling order of the modes is determined by imperfections and boundary conditions.
Matrix-Vector Multiplication – Computing in materia is an emerging direction [56, 32], studied in a myriad of physical platforms, such as cross-bar arrays [57], spintronic devices [58], electromagnetic [59, 60] and phononic coupled resonators [61], and microfluidics [62]. For these, the ability to perform algebraic operations is paramount, and therefore a crucial challenge is how to achieve matrix manipulations. Recent work has shown that floppy modes can be used to achieve matrix-vector multiplication in mechanical systems [35]. Here, we demonstrate that frustrated loops are an efficient way to perform algebraic operations with multiple inputs and outputs.
Figure 4: Matrix-vector multiplication. a, Experimental setup wherein the six corner nodes are fixed in place. Arrows indicate the decay in displacement along the two branches of the frustrated loop (dashed). b, Experimental results (dots) and theoretical predictions with $\alpha=0.8$ (solid line) of output under a single input from the top, $I_{\text{T}}$ or the bottom, $I_{\text{B}}$. Colors correspond to the outputs marked in panel a. c, Experimental measurements (solid lines) and predictions (dashed lines) under two simultaneous inputs. d, Three-layer metamaterial. e, Corresponding 3D structure. f, Experiment results (solid lines) of inputs in the middle layer, and outputs in the left and right layers. The dashed lines are the predictions.
The minimal metamaterial with a frustrated loop consists of six $T_{2}$ blocks (Fig. 4a). To enable manual actuation, we 3D print the diamond-shaped beams in this metamaterial with a flexible material and print the ligaments with a softer material. We fix the hexagon’s six corner nodes and locally actuate the top and bottom edge nodes, as input displacements $I_{\rm T}$ and $I_{\rm B}$. In the ideal model of rigid bonds, realized above using LEGO®, a chain with an odd loop is rigid, and may not deform. However, the 3D-printed metamaterial has some flexibility, leading to non-trivial output displacements $O_{\rm TL}$, $O_{\rm TR}$, $O_{\rm BL}$, and $O_{\rm BR}$ (Video 4). Actuating each of the inputs separately, we observe linear behavior of each of the outputs (Fig. 4b). Thus, this frustrated loop relates the input, $\vec{I}=\begin{pmatrix}I_{\rm T},I_{\rm B}\end{pmatrix}^{T}$, and output $\vec{O}=\begin{pmatrix}O_{\rm TL},O_{\rm TR},O_{\rm BL},O_{\rm BR}\end{pmatrix}^{T}$ vectors via the matrix-vector product, $\vec{O}=A\vec{I}$.
Each internal bond’s effective antiferromagnetic interaction flips the displacement’s direction. Due to the odd length of frustrated loops, the transmission of displacement from the input node along the loop’s two branches leads to displacement in opposite directions at the output node. However, beyond these flips in direction, the elastic nature of the 3D-printed metamaterial causes a decay in the displacement’s magnitudes [63], which we theoretically describe by each beam decreasing the displacement by some factor $\alpha$. Tracking the two paths from each input to each output node in the hexagonal structure of Fig. 4a, we obtain the transmission matrix,
$A$ $=\begin{pmatrix}-\alpha^{4}+\alpha^{7}&-\alpha^{2}+\alpha^{7}&\alpha^{4}-\alpha^{7}&\alpha^{4}-\alpha^{7}\\ -\alpha^{3}+\alpha^{8}&\alpha^{4}-\alpha^{5}&\alpha^{3}-\alpha^{8}&-\alpha^{4}+\alpha^{7}\end{pmatrix},$ (2)
which is consistent with the experimental measurement (Fig. 4b). Simultaneously applying two inputs in the same direction and then in opposite directions, we see that superposition holds (Fig. 4c), such that the output displacements are indeed given by $\vec{O}=A\vec{I}$.
Each frustrated loop may receive two inputs. Multiplying larger vectors is possible by feeding the output from one loop as an input to another loop. However, this requires transmission lines to cross, which is problematic in 2D metamaterials. We overcome this by coupling adjacent layers in a 3D metamaterial (Fig. 4d,e). Individually actuating each of the input nodes in the middle layer and measuring the response of each of the output nodes in the left and right layers, we obtain all elements of the transmission matrix $A$. We test our approach by simultaneously applying different combinations of several inputs (Fig. 4f), and observe that the output displacements are well described by $\vec{O}=A\vec{I}$.
We introduced a combinatorial approach using unimodal and bimodal triangular building blocks to create metamaterials with arbitrary shapes of multiple floppy modes and frustrated loops. We demonstrated that curved floppy modes together with yield buckling can be used to achieve sequential deformations, and that frustrated loops can be exploited to achieve matrix-vector multiplication. Our findings enrich the toolbox of sequential metamaterial design and lead to various applications in shock and vibration damping, shape changing, and mechanical computing. Our work opens questions like: can we use plasticity and fatigue to learn in combinatorial metamaterials [64]? How can frustrated loops integrate matrix-vector multiplication and nonlinear response like buckling for the function of logic gates [49]? How can floppy modes and frustrated loops be used for locomotion in active metamaterials [65]?
Acknowledgments – We thank Chaviva Sirote-Katz, Daniela Kraft, Dor Shohat, Julio Melio and Marc Serra-Garcia for inspiring discussions and Bat-El Pinchasik and Yaara Shokef for technical assistance. T.A.S. and Y.S. thank the Institute of Physics at the University of Amsterdam for its hospitality. T.A.S. acknowledges funding from the Israeli Ministry of Energy and Infrastructure’s scholarship program for undergraduate students in the field of energy. W.L. and C.C. acknowledge funding from the European Research Council under grant agreement 852587 and the Netherlands Organisation for Scientific Research under grant agreement NWO TTW 17883. C.C. acknowledges funding from the Netherlands Organisation for Scientific Research under grant agreement NWO under grant agreement VI.Vidi 2131313. This research was supported in part by the Israel Science Foundation Grant No. 1899/20.
Code and Data Availability – The supplementary videos as well as all the data and computer codes supporting this work are available at https://doi.org/10.5281/zenodo.15130951.
I Appendix
I.1 Bounds on Number of Floppy Modes
Figure 5: Average number of floppy modes in randomly generated lattices. a, Results for different system sizes normalized as $\tilde{F}=F/[(N+P)/2]$ and $\tilde{N_{1}}=[N_{1}-(N-P)/2]/[(N+P)/2]$, such that the lower bound will coincide. b, Normalized umber of floppy modes at the onset of rigidity, $\tilde{N}_{1}=0$.
To maximize the number $F$ of floppy mode, we minimize the number $R$ of rigid chains and maximize the number $L$ of loops. The minimal value of $R$ is zero. We now derive two upper bounds on $L$. The first stems from the fact that each loop surrounds at least one internal corner node, thus $L\leq\frac{N+2-P}{2}$, which leads to $F\leq N_{1}+1$.
For the second bound, we start with a lattice of $N_{2}=0$ and all triangles as $T_{1}$; In this limit, the maximal number of loops gives $L\leq\frac{N}{6}$, by arranging the loops in a honeycomb formation. Now, as we gradually switch $T_{1}$ blocks to $T_{2}$, the most efficient way to create new loops is that two loops share each $T_{2}$ block. Each new loop needs six bonds, thus $L\leq\frac{N}{6}+\frac{N_{2}}{3}$, and $F\leq\frac{2N_{1}}{3}+\frac{P}{2}$. The crossover between these two bounds occurs at $N_{1}=\frac{3P}{2}-3$, and we obtain the upper bound given in Eq. (1).
For a given lattice of size and shape, the system becomes more rigid as there are more $T_{2}$ than $T_{1}$ blocks. We obtain the minimal number $N_{2}$ of $T_{2}$ blocks or maximal number $N_{1}$ of $T_{1}$ blocks required for rigidity by requiring that all $E$ edge nodes belong to a single chain with at least one odd loop. Thus, the number of internal bonds must be at least $E$, with $E-1$ bonds connecting all nodes and one bond closing an odd loop. Thus, mechanical rigidity can be achieved when $N_{1}\leq\frac{N-P}{2}$. In this case $F=0$. As we increase $N_{1}$, we remove internal bonds. There are no loops, and this removal of bonds does not create loops. Thus, by substituting in the expression for the number of modes, we obtain the lower bound given in Eq. (1).
In the large system size limit, $N\gg 1$, since $P\ll N$, the lower bound is approximated by
$F\geq\begin{cases}0,&N_{1}\leq\frac{N}{2}\\ N_{1}-\frac{N}{2},&N_{1}\geq\frac{N}{2},\end{cases}$ (3)
and the upper bound by $F\leq\frac{2N_{1}}{3}$. As system size increases, the average number of floppy modes in randomly generated lattices collapses to a single curve (Fig. 5).
I.2 Multi-Layer Matrix-Vector Multiplication
We have demonstrated the matrix-vector multiplication in a single-layer metamaterial. Here, we further show a larger matrix-vector multiplication in a three-layer metamaterial of four inputs in the middle layer and 12 outputs in the side layers. These frustrated loops similarly relate the input and output vectors. We experimentally apply a single input and measure the transmission matrix $A$ from the slopes of the input-output curves (Fig. 6). Similarly to the single-layer case, we predict the transmission by assuming decay factor $\alpha=0.8$ within each layer, decay factor $\beta_{1}=0.4$ when the input transmits from one layer to another layer, and $\beta_{2}=1$ when two layers have input and output at the same location, like the output $L_{\rm T}$ for the input $I_{\rm T}$ (Fig. 4d). We find good agreement between the theoretical predictions and the experimental measurements (Fig. 6).
Figure 6: Experimental outputs of the three-layer metamaterial with single inputs. a, Left layer. b, Right layer. Inputs are applied to the middle layer. Solid lines are the theoretical predictions.
Figure 7: 3D printed metamaterials and experimental method. a, Metamaterial for buckling with elastoplastic material (Nylon). b, Metamaterial with tracking markers at the connected joints. c, Stretching stress-strain curve of 3D printed Nylon where we linear fit the elastic modulus and tangent modulus. The white stickers on the dogbone sample are used to track the strain. d, A single-layer metamaterial for local actuation is printed with two materials. Purple ligaments and red diamonds are printed with soft material (Agilus) and rigid material (Vero), respectively. e, Experimental setup of a three-layer metamaterial with local actuation under filming of two cameras. The picture in the top-right corner is the recorded image from the right camera
I.3 Experimental Methods
The metamaterials with floppy modes designed in this work consist of a plurality of triangular unit cells, wherein the rigid bonds and rotating nodes from the theoretical model are replaced with solid diamond blocks of length $\ell$ and width $d$ and ligaments of thickness $t$, respectively. For the metamaterials with buckling floppy modes, we 3D printed them with a single material of $\ell=10.3$ mm, $t=0.4$ mm, and $d=3$ mm (Fig. 7a). For the metamaterials with yield buckling, we used an elastoplastic material, nylon (with a low ratio of $E_{t}/E=0.56\%$ see calibration in Fig. 7c), with the fused deposition modelling (FDM) technology (Ultimaker 2S). For the metamaterials with elastic buckling, we 3D printed them with a rubber-like polymer material (Stratasys, Agilus+Vero). The thicknesses of the metamaterials are set to $D=24.5$ mm for elastoplastic metamaterials and $D=23.2$ mm for elastic metamaterials to prevent out-of-plane buckling. We then marked the ligaments on one side of the metamaterials with white dots for image tracking (Fig. 7b). We used Image-J software to track the position of the white dots of each frame in the video. For the metamaterials with local actuation, we changed the design of ligaments from a rectangular section to a section with inserted tapered rigid parts (Fig. 7d). We then printed such ligaments with a soft material (Agilus) and the diamonds with a rigid material (Vero). Such ligament design can significantly increase resistance to shear and tension while offering only a negligible resistance to bending [39]. For the single-layer metamaterial (Fig. 4b), we set the thickness of the sample to 6 mm. For the three-layer metamaterials with multiple inputs and outputs, we printed each layer separately and assembled them with two rigid frames. The middle layer has a thickness of 6 mm and the left and right layers have a thickness of 3.6 mm. We set a small distance of 0.5 mm between each layer to avoid friction during actuation. We then fixed the sample on the table. We marked the ligaments at the six edges of the hexagon layer with black dots (Fig. 7e top-right corner). We then manually pulled and pushed the rigid bars of the middle layer and recorded the movement of the 12 dots from the left and right layers with two Basler cameras (Fig. 7e).